OpenStudy (anonymous):

Round the fraction 33 55/68

OpenStudy (michele_laino):

we can write this: \[33\,{\kern 1pt} {\kern 1pt} \frac{{55}}{{68}} = \frac{{33 \cdot 68 + 55}}{{68}} = ...?\]

OpenStudy (anonymous):

33

OpenStudy (michele_laino):

how many decimal figures do you need?

OpenStudy (anonymous):

decimal figures?

OpenStudy (michele_laino):

since we have: \[33\,{\kern 1pt} {\kern 1pt} \frac{{55}}{{68}} = \frac{{33 \cdot 68 + 55}}{{68}} = 33.80882353...\] if one decimalfigure is requested, the the rounded off form is: \(33.8\)

OpenStudy (anonymous):

So it would be 35

OpenStudy (michele_laino):

if none decimal figure is requested, then the rounded off form is \(34\)

OpenStudy (anonymous):

Thank you :)

OpenStudy (michele_laino):

:)